Thanks in advance!

Constants | Periodic Table Two point charges lie on the x-axis: 1.65 pC at z = 1.00 m and-220 pC at z=2.55 m .
Find the place(s) where the electric field is zero. Express your answer using two decimal places. If there is more than one answer, separate them by commas. In Submit
Answer #1

Here,

let electric field is zero left of 1.65 uC charge ,

let the distance left of 1.65 uC is d

Now, for the electric field

k * 1.65/d^2 = k * 2.2/(2.55 -1 + d)^2

1.65/d^2 = 2.2/(2.55 -1 + d)^2

solving for d

d = 10.02 m

x position = 1 - 10.02 = -9.02 m

the position where electric field is zero is -9.02 m